# Mellivora Capensis: A Backdoor-Free Training Framework on the Poisoned Dataset without Auxiliary Data

### v2

@@ Abstract
The efficacy of deep learning models is profoundly influenced by the quality of their training data. Given the considerations of data diversity, data scale, and annotation expenses, model trainers frequently resort to sourcing and acquiring datasets from online repositories. Although economically pragmatic, this strategy exposes the models to substantial security vulnerabilities. Untrusted entities can clandestinely embed triggers within the dataset, facilitating the hijacking of the trained model on the poisoned dataset through backdoor attacks, which constitutes a grave security concern. Despite the proliferation of countermeasure research, their inherent limitations constrain their effectiveness in practical applications. These include the requirement for substantial quantities of clean samples, inconsistent defense performance across varying attack scenarios, and inadequate resilience against adaptive attacks, among others. Therefore, in this paper, we endeavor to address the challenges of backdoor attack countermeasures in real-world scenarios, thereby fortifying the security of training paradigm under the data-collection manner. Concretely, we first explore the inherent relationship between the potential perturbations and the backdoor trigger, and demonstrate the key observation that the poisoned samples perform more robustness to perturbation than the clean ones through the theoretical analysis and experiments. Then, based on our key explorations, we propose a robust and clean-data-free backdoor defense framework, namely Mellivora Capensis (MeCa), which enables the model trainer to train a clean model on the poisoned dataset. MeCa can effectively identify poisoned samples and train a clean model on the poisoned dataset and eliminate the requirement of clean samples or knowledge about the poisoned dataset (e.g., poisoning ratio). We conduct extensive experiments in defending against $8$ SOTA attacks (including $3$ adaptive attacks) on $4$ datasets (Imagnette, Tiny ImageNet, CIFAR10, and CIFAR100). The experimental results reveals that MeCa can achieve an average ASR with almost $0.00\%$ to defend against SOTA backdoor attacks while maintaining model availability, which outperforms $7$ SOTA backdoor defense methods. Furthermore, the excellent performance on $3$ different model architectures and poisoning ratios also highlights the remarkable generalization capability of MeCa.
Index Terms:
backdoor attack, backdoor defense, adversarial perturbation, model security.
1 Introduction
@@
Training deep learning models relies on a large amount of training data [1, 2, 3, 4]. However, the training data are usually collected from the web or purchased from an untrustworthy third party, which leads to a risk of being poisoned or unreliable [5, 6]. For example, a company that wants to train a facial recognition model is likely to obtain poisoned data from a third party who aims to hijack the facial recognition model for profit [7, 8], as depicted in Fig.1. Therefore, the unreliable training dataset may bring many serious security threats to the model training side, which breaches the national regulations, such as Artificial Intelligence Act (AIA) [9], Digital Services Act (DSA) [10]. One of the typical threats is the backdoor attack, which is usually conducted by poisoning the training data [11, 12]. In the backdoor attack, the attacker manipulates a few training samples by adding a specific trigger (e.g., a pixel patch or blending) and modifies the labels of these samples as a target class. The backdoor model maintains high accuracy on the normal samples but predicts the target class whenever the trigger pattern is attached to a test sample, as depicted in Fig.1. Backdoor attacks bring a significant security risk for model training owners as they are easily conducted and allow adversaries to control the trained model stealthily [13, 14]. For example, a backdoored traffic sign recognition model may misclassify a stop sign with the trigger patch as a no-tooting sign, which may cause a serious traffic accident. Therefore, untrusted sources of datasets often serve as a breeding ground for backdoor attacks, posing a significant security threat to model trainers.
@@
Regarding the third-party poisoned training dataset, both academia and industry have conducted extensive research, such as the detection and purification of backdoor samples in poisoned datasets [15, 16]. Although effective backdoor detection and purification are valuable, economically speaking, it is clearly more cost-effective and practical to train a clean model directly on a poisoned dataset if that were possible. Such paradigm simplifies data prep and reduces cleaning costs by enabling training on potentially tainted datasets, fostering robust models without extensive preprocessing. It’s crucial in resource-intensive data scenarios and could redefine data handling across fields, ensuring reliable, cost-effective outcomes [17], which implies we possess the capability to rebuild amidst the ruins.
@@
Consequently, another research direction in backdoor defense is to train clean models on poisoned datasets [18, 19, 20, 21, 22]. However, the existing methods have some non-negligible limitations in practical applications. (1) Requirement for clean samples. Their defense effectiveness is highly dependent on the obtaining of a substantial number of clean samples [20, 21, 22], which is impractical for defenders in certain specific real-world scenarios [23]. (2) Unstable defense performance. Their actual defense performance varies significantly depending on the choice of backdoor attack methods and poisoning ratios [24, 19], which is usually unknown to defenders in advance in practical applications. (3) Inadequate defense against adaptive attacks. They may exhibit poor resilience against adaptive attacks [19], e.g., transforming some poisoned samples.
Therefore, in this context, we ask the following research questions: {mdframed}[backgroundcolor=black!10,rightline=false,leftline=false,topline=false,bottomline=false,roundcorner=2mm] Is it feasible to train a clean model directly on a poisoned dataset, without any knowledge of the attacker’s capabilities (such as poisoning ratio) and without any auxiliary clean dataset? If so, how to ensure model availability?
@@
  • How to identify the poisoned samples regardless of the poisoning ratio and without the auxiliary clean dataset?
@@
  • We explore the intrinsic differences between poisoned samples and clean samples, uncovering key observation about the differences in their perturbation manifestations. We conducted extensive theoretical analysis and experiments to substantiate our findings.
@@
  • We evaluate MeCa through extensive experiments on four datasets, which demonstrate a better backdoor defense performance compared to seven SOTA backdoor defense methods. Experimental results validate that MeCa could decrease the attack success rate (average ASR nearly 0.00%) while only incurring negligible accuracy loss. Furthermore, MeCa also demonstrates robust generalization across various model architectures and poisoning ratios.
2 Related Works
@@
2.1 Backdoor Attack Methods
@@
In recent years, many backdoor attack approaches have been proposed, including patch-based attacks, visibility of trigger attacks, label consistency, and so on. Gu et al.[12] proposed a typical backdoor attack method by adding a specific patch on the samples and modifying the corresponding label as the targeted label. This attack achieves high attack success and has a low-performance impact on the main task. Chen et al.[25] proposed a backdoor attack in which the trigger is blending background rather than a single pixel. This means that the trigger is hard for human beings to notice. Liu et al.[26] proposed a more stealthy backdoor attack that plants reflections as a trigger into the victim model. This attack can be resistant to many existing defense methods. Liu et al.[27] proposed a lightweight backdoor attack that just needs to inverse the neural network to generate a general trojan trigger and fine-tune some layers of the network to implant the trigger. However, all the above backdoor attacks require poisoning the label, which is easier to detect. Therefore, some researchers have also proposed clean-label backdoor attacks. For example, Shafahi et al.[28] proposed a targeted clean-label poisoning attack. This attack crafts poison images that collide with a target image in feature space, thus making it undistinguishable from a network. Because the attacker does not need to control the label, it is more stealthy to conduct a backdoor attack. Turner et al.[29] proposed a clean-label backdoor attack based on adversarial examples and GAN-generated data. The key feature of this attack is that the poisoned samples appear to be consistent with their label and thus seem clean even from human inspection. Chen et al.[8] proposed an Invisible Poisoning Attack (IPA), which is difficult to detect by existing defense methods. This attack not only employs highly stealthy poison training examples with the clean labels (perceptually similar to their clean samples), but also does not need to modify the labels. Li et al.[30] proposed two stealthy backdoor attacks in which the triggers are derived from the covert features. Compared with the existing backdoor attacks, the trigger patterns of this attack method are invisible to human eyes. Moreover, it is difficult to recover the backdoor trigger through the optimization algorithm. Zhu et al.[31] proposed a transferable clean-label poisoning attack in which poison samples are fabricated to surround the targeted sample in feature space. Saha et al.[32] proposed a novel backdoor attack in which the poisoned samples are similar to the clean samples with the correct labels. [33] and [34] proposed defense-resistant backdoor attacks in an outsourced cloud environment.
2.2 Backdoor Defense Methods
@@
Many defense methods are proposed to resist the existing backdoor attacks. Some approaches require that the defender must own some reserved clean datasets. For example, Qi et al.[20] proposed a backdoor sample detection method that directly enforces and magnifies distinctive characteristics of the post-attacked model to facilitate poison detection. Guo et al.[21] proposed a universal detection approach based on clustering and centroids analysis. The approach can detect the poisoned samples based on density-based clustering and the clean validation dataset. Zhu et al.[35] proposed a defense method by inserting a learnable neural polarizer layer, which is optimized based on a limited clean dataset. The layer can purify the poisoned sample by filtering trigger information while maintaining clean information. For a backdoored model, Chen et al.[22] proposed a generic scheme for defending against backdoor attacks. The insight of this scheme is to localize the neuron set related to the trigger with the auxiliary clean dataset and suppress the compromised neurons. Ma et al.[36] proposed an input-level detection method. The intuition of this method is that even though a poisoned sample and a clean sample are classified into the target label, their intermediate representations are also different. Based on this observation, the poisoned samples can be detected easily. Wei et al.[37] presented a backdoor mitigation method using a small clean dataset. This method employs unlearning shared adversarial examples to purify the backdoored model. Researchers also proposed some backdoor defense methods without auxiliary clean datasets. It is a more strict and practical setting. For instance, Li et al.[18] proposed a defense approach that aims to train the clean model on the poisoned data. The main intuition of this method is that the models learn backdoor samples much faster than learning with clean samples. The backdoor examples can be easily removed by filtering out the low-loss examples at an early stage. Because the poisoned samples are much more sensitive to transformations than the clean samples in a backdoored model, Chen et al.[19] distinguish the poisoned samples from clean samples based on the feature consistency towards transformations. Weng et al.[38] found a trade-off between adversarial and backdoor robustness. Then, Gao et al.[39] challenged the trade-off between adversarial and backdoor robustness and proposed a backdoor defense strategy based on adversarial training regardless of the trigger pattern. Li et al.[40] found that the existing backdoor attacks have non-transferability. That is, the trigger sample is not effective in another model that has not been injected with the same trigger. Based on this observation, the authors proposed an input sample detection method by comparing the input sample and the samples picked from its predicted class label based on a feature extractor. Huang et al.[41] proposed a backdoor defense via decoupling the training process, thereby breaking the connection between the trigger and target label. Mu et al.[42] observed that the adversarial examples have similar behaviors as the triggered samples. Then, a progressive backdoor erasing method is proposed to purify the poisoned model via employing untargeted adversarial attacks. Tang et al.[43] presented a robust backdoor detection approach that can effectively detect data contamination attacks. Feng et al.[44] proposed a backdoor detection method for pre-trained encoders, requiring neither classifier headers nor input labels. Chen et al.[45] presented a robust backdoor defense scheme for federated learning. This scheme can overcome many backdoor attacks, including amplified magnitude sparsification, adaptive clipping, and so on.
@@
Although many backdoor defense methods have been proposed, they have some limitations (e.g., requiring an auxiliary clean dataset) used in practical scenarios. Unlike existing backdoor defense methods, we first explore the relationship between the perturbation and the backdoor. Then, based on the exploration results, we plan to propose a robust backdoor defense method that does not require an auxiliary clean dataset and has a stable performance against various backdoor attacks on different models and poisoning ratios.
3 Preliminaries and Threat Model
@@
In this section, we introduce the mainly related technologies and the threat model in this paper.
3.1 Backdoor Attack
@@
A backdoor attack is to inject a trigger into a model by poisoning the training dataset. During the inference period, the backdoored model performs well on the original task, but outputs specific attacker-chosen responses when the input contains a specific trigger. For more clarity, we formalize the most common backdoor attack method BadNets [12] as follows:
@@
Let $f:\mathcal{X}\rightarrow\mathcal{Y}$ be a neural network for an image classification task. Let $\mathcal{D}=\left\{\left(\boldsymbol{x}_{i},y_{i}\right)\right\}_{i=1}^{N}$ be a clean training set where $\boldsymbol{x_{i}}\in\mathcal{X}$ and $\boldsymbol{y_{i}}\in\mathcal{Y}$ are the training image and the corresponding label, respectively. To conduct a backdoor attack, the attacker will choose a backdoor trigger $\boldsymbol{b}=(\boldsymbol{m},\boldsymbol{t})$ that consists of a blending mask $\boldsymbol{m}$ and a trigger pattern $\boldsymbol{t}$. In general, the trigger pattern is usually small just to achieve a stealthier backdoor attack. During the training process, the attacker randomly selects some training samples and poisons them by adding a specific backdoor trigger. For one poisoned training sample:
$\boldsymbol{x}^{\prime}=(\mathbf{1}-\boldsymbol{m})\times\boldsymbol{x}+\boldsymbol{m}\times\boldsymbol{t}.$ (1)
@@
where $\times$ is pixel-wise multiplication and $\boldsymbol{x}^{\prime}$ is the poisoned training sample. After modifying the training sample, the corresponding label is fixed with a target label. A successful backdoor model will maintain high accuracy in the main task while outputting the target label when the trigger $\boldsymbol{t}$ appears. After BadNets [12], many different attack variations have been proposed to enhance effectiveness and stealthiness. For example, Blend [25] and PhysicalBA [46] have designed more complex patterns. WaNet [47] has proposed a stealthier input-specific-trigger attack. SIG [48] has proposed a clean-label attack that is stealthier as it does not change the labels. To evade the existing defense method, some adaptive attacks (e.g., TaCT [43], AdaptiveBlend [49] and AdaptivePatch [49]) have also been proposed to improve the attack performance.
@@
3.2 Threat Model
In the scenario considered in MeCa, the model trainer (Defender) intends to train a model on the training data obtained from a third party. A malicious data provider (Attacker) endeavors to manipulate the dataset in order to induce the defender to train a model with a backdoor, thereby facilitating subsequent hijacking of the model. The threat model is depcited in Fig.2. Therefore, in this paper, the defender utilizes MeCa to train a clean model on the poisoned dataset, free from the attacker’s backdoor triggers. Considering the deployment and development of MeCa in real-world scenarios, we provide a detailed description of the goals, knowledge, and capabilities of both the Defender and the Attacker in the following.
Defender’s goal: The defender aims to identify the poisoned samples and train a clean model on the poisoned dataset.
Defender’s knowledge and capabilities: The defender defines the model architecture, learning algorithm, and hyperparameters and trains the model. It has no knowledge about the poisoning ratio and without any auxiliary clean dataset.
Attacker’s goal: The attacker aims to inject a backdoor into the trained model by poisoning the training dataset.

### v3

@@ Abstract
Deep learning models heavily depend on training data quality. While online datasets offer cost-effective solutions for diversity and scale, they introduce security risks. Malicious actors can inject hidden triggers, enabling backdoor attacks that compromise model integrity. Existing defenses remain limited—often demanding large clean datasets, showing inconsistent robustness across attacks, and struggling against adaptive adversaries. Therefore, in this paper, we endeavor to address the challenges of backdoor attack countermeasures in real-world scenarios, thereby fortifying the security of the training paradigm under the data-collection manner. Concretely, we first explore the inherent relationship between the robustness of the poisoned samples, demonstrating the poisoned samples are more robust to perturbation than the clean ones through the theoretical analysis and experiments. Then, we propose a robust and clean-data-free backdoor defense framework, namely Mellivora Capensis (MeCa), which enables training a clean model on the poisoned dataset. MeCa detects poisoned samples and trains clean models without needing clean data or prior knowledge of the poisoning (e.g., poison ratio). We conduct extensive experiments in defending against $8$ SOTA attacks (including $3$ adaptive attacks) on $4$ datasets. The experimental results reveal that MeCa can achieve an average attack success rate with almost $0.00\%$ to defend against SOTA backdoor attacks while maintaining model availability, which outperforms $7$ SOTA backdoor defense methods. Furthermore, the excellent performance on $3$ different model architectures and poison ratios also highlights the remarkable generalization capability of MeCa.
I Introduction
@@
Training deep learning models relies on a large amount of training data [1, 2, 3, 4]. However, the training data are usually collected from the web or purchased from an untrustworthy third party, which leads to a risk of being poisoned or unreliable [5, 6]. For example, a company that wants to train a facial recognition model is likely to obtain poisoned data from a third party who aims to hijack the facial recognition model for profit [7, 8, 9], as depicted in Fig.1. Therefore, the unreliable training dataset may bring many serious security threats to the model training side, which breaches the national regulations, such as Artificial Intelligence Act (AIA) [10], Digital Services Act (DSA) [11]. One of the typical threats is the backdoor attack, which is usually conducted by poisoning the training data [12, 13]. In the backdoor attack, the attacker manipulates a few training samples by adding a specific trigger (e.g., a pixel patch or blending) and modifies the labels of these samples as a target class. The backdoor model maintains high accuracy (ACC) on the normal samples but predicts the target class whenever the trigger pattern is attached to a test sample, as depicted in Fig.1. Backdoor attacks bring a significant security risk for model training owners as they are easily conducted and allow adversaries to control the trained model stealthily [14, 15, 16, 17, 18, 19]. For example, a backdoored traffic sign recognition model may misclassify a stop sign with the trigger patch as a no-tooting sign, which may cause a serious traffic accident. Therefore, untrusted sources of datasets often serve as a breeding ground for backdoor attacks, posing a significant security threat to model trainers.
@@
Regarding the third-party poisoned training dataset, both academia and industry have conducted extensive research, such as the detection and purification of backdoor samples in poisoned datasets [20, 21, 22]. Although effective backdoor detection and purification are valuable, economically speaking, it is clearly more cost-effective and practical to train a clean model directly on a poisoned dataset if that were possible. Such a paradigm simplifies data preparation and reduces cleaning costs by enabling training on potentially tainted datasets, thereby fostering robust models without the need for extensive preprocessing. A recent study [23] also validated that it’s hard to distinguish a poisoned sample from a clean one even for human experts. Therefore, developing a data-free robust training strategy in resource-intensive data scenarios could redefine data handling across fields, ensuring reliable, cost-effective outcomes [24], which implies we possess the capability to rebuild amidst the ruins.
@@
Consequently, another research direction in backdoor defense is to train clean models on poisoned datasets [25, 26, 27, 28, 29]. However, the existing methods have some non-negligible limitations in practical applications. (1) Requirement for clean samples. Their defense effectiveness is highly dependent on the obtaining of a substantial number of clean samples [27, 28, 29], which is impractical for defenders in certain specific real-world scenarios [30]. (2) Unstable defense performance. Their actual defense performance varies significantly depending on the choice of backdoor attack methods and poison ratios [31, 26], which is usually unknown to defenders in advance in practical applications. (3) Inadequate defense against adaptive attacks. They may exhibit poor resilience against adaptive attacks [26], e.g., transforming some poisoned samples.
Therefore, in this context, we ask the following research questions: Is it feasible to train a clean model directly on a poisoned dataset, without any knowledge of the attacker’s capabilities (such as poison ratio) and without any auxiliary clean dataset? If so, how to ensure model availability?
@@
  • How to identify the poisoned samples regardless of the poison ratio and without the auxiliary clean dataset?
@@
  • We explore the intrinsic differences between poisoned samples and clean samples, uncovering key observation about the differences in their perturbation manifestations. We conduct extensive theoretical analysis and experiments to substantiate our findings.
@@
  • We evaluate MeCa through extensive experiments on four datasets, which demonstrate a better backdoor defense performance compared to seven SOTA backdoor defense methods. Experimental results validate that MeCa could decrease the attack success rate (ASR) (nearly 0.00%) while only incurring negligible accuracy loss. Furthermore, MeCa also demonstrates robust generalization across various model architectures and poison ratios.
II Related Works
@@
II-A Backdoor Attack Methods
@@
In recent years, many backdoor attack approaches have been proposed, including patch-based attacks, visibility of trigger attacks, label consistency, and so on. Gu et al.[13] proposed a typical backdoor attack method by adding a specific patch on the samples and modifying the corresponding label as the targeted label. This attack achieves high attack success and has a low-performance impact on the main task. Chen et al.[32] proposed a backdoor attack in which the trigger is blending background rather than a single pixel. This means that the trigger is hard for human beings to notice. Liu et al.[33] proposed a stealthier backdoor attack that plants reflections as a trigger into the victim model. This attack can be resistant to many existing defense methods. Liu et al.[34] proposed a lightweight backdoor attack that just needs to inverse the neural network to generate a general trojan trigger and fine-tune some layers of the network to implant the trigger. However, all the above backdoor attacks require poisoning the label, which is easier to detect. Therefore, some researchers have also proposed clean-label backdoor attacks. For example, Shafahi et al.[35] proposed a targeted clean-label poisoning attack. This attack crafts poison images that collide with a target image in feature space, thus making it undistinguishable from a network. Because the attacker does not need to control the label, it is more stealthy to conduct a backdoor attack. Turner et al.[36] proposed a clean-label backdoor attack based on adversarial examples and GAN-generated data. The key feature of this attack is that the poisoned samples appear to be consistent with their label and thus seem clean even from human inspection. Chen et al.[8] proposed an Invisible Poisoning Attack (IPA), which is difficult to detect by existing defense methods. This attack not only employs highly stealthy poison training examples with the clean labels (perceptually similar to their clean samples), but also does not need to modify the labels. Li et al.[37] proposed two stealthy backdoor attacks in which the triggers are derived from the covert features. Compared with the existing backdoor attacks, the trigger patterns of this attack method are invisible to human eyes. Moreover, it is difficult to recover the backdoor trigger through the optimization algorithm. Zhu et al.[38] proposed a transferable clean-label poisoning attack in which poison samples are fabricated to surround the targeted sample in feature space. Saha et al.[39] proposed a novel backdoor attack in which the poisoned samples are similar to the clean samples with the correct labels. [40] and [41] proposed defense-resistant backdoor attacks in an outsourced cloud environment.
II-B Backdoor Defense Methods
@@
Many defense methods [42, 43, 44] are proposed to resist the existing backdoor attacks. Some approaches require that the defender must own some reserved clean datasets. For example, Qi et al.[27] proposed a backdoor sample detection method that directly enforces and magnifies distinctive characteristics of the post-attacked model to facilitate poison detection. Guo et al.[28] proposed a universal detection approach based on clustering and centroids analysis. The approach can detect the poisoned samples based on density-based clustering and the clean validation dataset. Zhu et al.[45] proposed a defense method by inserting a learnable neural polarizer layer, which is optimized based on a limited clean dataset. The layer can purify the poisoned sample by filtering trigger information while maintaining clean information. For a backdoored model, Chen et al.[29] proposed a generic scheme for defending against backdoor attacks. The insight of this scheme is to localize the neuron set related to the trigger with the auxiliary clean dataset and suppress the compromised neurons. Ma et al.[46] proposed an input-level detection method. The intuition of this method is that even though a poisoned sample and a clean sample are classified into the target label, their intermediate representations are also different. Based on this observation, the poisoned samples can be detected easily. Wei et al.[47] presented a backdoor mitigation method using a small clean dataset. This method employs unlearning shared adversarial examples to purify the backdoored model. Researchers also proposed some backdoor defense methods without auxiliary clean datasets. It is a more strict and practical setting. For instance, Li et al.[25] proposed a defense approach that aims to train the clean model on the poisoned data. The main intuition of this method is that the models learn backdoor samples much faster than learning with clean samples. The backdoor examples can be easily removed by filtering out the low-loss examples at an early stage. Because the poisoned samples are much more sensitive to transformations than the clean samples in a backdoored model, Chen et al.[26] distinguish the poisoned samples from clean samples based on the feature consistency towards transformations. Weng et al.[48] found a trade-off between adversarial and backdoor robustness. Then, Gao et al.[49] challenged the trade-off between adversarial and backdoor robustness and proposed a backdoor defense strategy based on adversarial training regardless of the trigger pattern. Li et al.[50] found that the existing backdoor attacks have non-transferability. That is, the trigger sample is not effective in another model that has not been injected with the same trigger. Based on this observation, the authors proposed an input sample detection method by comparing the input sample and the samples picked from its predicted class label based on a feature extractor. Huang et al.[51] proposed a backdoor defense via decoupling the training process, thereby breaking the connection between the trigger and target label. Mu et al.[52] observed that the adversarial examples have similar behaviors as the triggered samples. Then, a progressive backdoor erasing method is proposed to purify the poisoned model via employing untargeted adversarial attacks. Tang et al.[53] presented a robust backdoor detection approach that can effectively detect data contamination attacks. Feng et al.[54] proposed a backdoor detection method for pre-trained encoders, requiring neither classifier headers nor input labels. Chen et al.[55] presented a robust backdoor defense scheme for federated learning. This scheme can overcome many backdoor attacks, including amplified magnitude sparsification, adaptive clipping, and so on.
@@
Although many backdoor defense methods have been proposed, they have some limitations (e.g., requiring an auxiliary clean dataset) used in practical scenarios. Unlike existing backdoor defense methods, we first explore the relationship between the perturbation and the backdoor. Then, based on the exploration results, we plan to propose a robust backdoor defense method that does not require an auxiliary clean dataset and has a stable performance against various backdoor attacks on different models and poison ratios.
III Preliminaries and Threat Model
@@
In this section, we introduce the main related technologies and the threat model in this paper.
III-A Backdoor Attack
@@
A backdoor attack injects a trigger into the model by poisoning the training dataset. During the inference period, the backdoored model performs well on the original task, but outputs specific attacker-chosen responses when the input contains a specific trigger. For more clarity, we formalize the most common backdoor attack method BadNets [13], below.
@@
Let $f:\mathcal{X}\rightarrow\mathcal{Y}$ be a neural network for an image classification task. Let $\mathcal{D}=\left\{\left(\boldsymbol{x}_{i},y_{i}\right)\right\}_{i=1}^{N}$ be a clean training set where $\boldsymbol{x_{i}}\in\mathcal{X}$ and $\boldsymbol{y_{i}}\in\mathcal{Y}$ are the training image and the corresponding label, respectively. To conduct a backdoor attack, the attacker will choose a backdoor trigger that consists of a blending mask $\boldsymbol{m}$ and a trigger pattern $\boldsymbol{t}$. In general, the trigger pattern is usually small just to achieve a stealthier attack. During the training process, the attacker randomly selects some training samples and poisons them by adding a specific backdoor trigger. For one poisoned training sample:
$\boldsymbol{x}^{\prime}=(\mathbf{1}-\boldsymbol{m})\odot\boldsymbol{x}+\boldsymbol{m}\odot\boldsymbol{t}.$ (1)
@@
where $\odot$ is element-wise multiplication and $\boldsymbol{x}^{\prime}$ is the poisoned training sample. After modifying the training sample, the corresponding label is fixed with a target label. A successful backdoor model will maintain high accuracy in the main task while outputting the target label when the trigger $\boldsymbol{t}$ appears. After BadNets [13], many different attack variations have been proposed to enhance effectiveness and stealthiness. For example, Blend [32] and PhysicalBA [56] have designed more complex patterns. WaNet [57] has proposed a stealthier input-specific-trigger attack. SIG [58] has proposed a clean-label attack that is stealthier as it does not change the labels. To evade the existing defense method, some adaptive attacks (e.g., TaCT [53], AdaptiveBlend [59] and AdaptivePatch [59]) have also been proposed to improve the attack performance.
@@
III-B Threat Model
In the scenario considered in MeCa, the model trainer (Defender) intends to train a backdoor-robust model on the untrusted training data obtained from a third party. A malicious data provider (Attacker) endeavors to manipulate the dataset in order to induce the defender to train a model with a backdoor, thereby facilitating subsequent hijacking (e.g., targeted misclassification or creating authentication bypasses) of the model. The threat model is depicted in Fig.2. Therefore, in this paper, the defender utilizes MeCa to train a robust model on the untrusted dataset, free from the attacker’s knowledge (e.g., triggers or target class). Considering the deployment and development of MeCa in real-world scenarios, we provide a detailed description of the goals, knowledge, and capabilities of both the Defender and the Attacker in the following.
Defender’s goal: The defender seeks to detect poisoned samples and train a robust model from untrusted data, maintaining high task accuracy while resisting trigger-based attacks.
Defender’s knowledge and capabilities: The defender defines the model architecture, learning algorithm, and hyperparameters and trains the model. It has no knowledge about the poison ratio, and without any auxiliary clean dataset, which is the recognized practice in real-world scenarios [23] since poisoned samples can hardly be distinguished even for human experts.
Attacker’s goal: The attacker compromises the model by injecting malicious samples into the training dataset to establish a backdoor.